# BINDy – Bayesian identification of nonlinear dynamics with reversible-jump Markov-chain Monte-Carlo

### v1

@@ 1 Introduction
Bayesian uncertainty quantification is a natural framework for conducting UQ in data-driven modelling. Indeed, much has been written on the topic, e.g. [2, 3]. Physical knowledge and inductive biases (such as parsimony) can be incorporated as prior knowledge, making such modelling assumptions explicit. Indeed, many literature contributions propose Bayesian methods for UQ in both model parameterisation [4, 5, 6] and model structure [7, 8, 9, 10]. Further discussion of specific related works is presended in Section 3.4.
@@ 1.1 Contribution
The remainder of this paper is structured as follows: The following section summarises the SINDy method and holds some discussion on parsimony in data-driven modelling. A third section reviews the reversible-jump theory and introduces the proposed approach and its position in the literature. A fourth section compares the proposed approach to the ensemble SINDy method on two benchmark case studies. A final section presents some discussion and directions for future work.
2 SINDy methods
The SINDy method proposed in 2016 by Brunton et al. [1] has received considerable attention in the scientific literature as a computationally-efficient way to learn differential equations directly from data. Since it was first proposed, the method has been proven to be effective in correctly identifying governing equations from both simulated and real-world data.
At their core, SINDy methods make the key assumption that the observed dynamics admit a first-order formulation that is linear in the parameters of a library of user-specified basis functions $\Theta$. Let $f$ denote the first-order ordinary differential equation (ODE) that describes the evolution of some state-space dynamics $x(t)$, [^1] The method can trivially be extended to systems of ODEs, PDEs etc. but for notational simplicity, the exposition in this paper will consider scalar $x$ and $f$ only.
$\dot{x}(t)=f(x(t))$ (1)
@@ 2 SINDy methods
$\dot{x}\approx\Theta(x)\Xi$ (2)
where $x$ is the state vector, $\Theta$ is a library of user-selected basis functions computed from $x$. Critically, the coefficient vector $\Xi$ is assumed to be sparse, whereby many elements of $\Xi$ are equal to 0. In practice, it is not often the case that measurements of both $x$ and $\dot{x}$ are available. It is usual in the SINDy literature to assume that only the states $x$ are measured and their derivatives $\dot{x}$ can be computed numerically (with sufficient accuracy such that (2) holds). The SINDy framework (for identification of ODEs) can thus be summarised by three steps:
@@
3 Proposed approach
Of the three components in SINDy methods described in the previous section, this paper is concerned only with the third—determination of appropriate model terms and their parameters. At this stage, it is useful to present some discussion as to the role of parsimony in data-driven modelling.
@@ 3.1 Parsimony in data-driven modelling
A particular issue with parameter size as a proxy for term importance is the selection of a threshold value, below which model terms should be excluded. In practice, this level can be difficult to select in a principled manner, leading to a spectrum of models at different levels of sparsity.
@@
$\dot{x}=\Theta(x)\Xi+\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(0,\sigma^{2})$ (3)
where $\sigma^{2}$ is the noise variance in the observation model. The model identification task at hand is to identify a subset of the columns of $\Theta(x)$. Intuitively, this model set includes all subsets of library functions (the powerset) and thus all possible ways that the parameter vector may be sparsified. Let,
=$\mathcal{M}=\{m_{i}\}_{i=1}^{2^{n}}$ (4)
@@
$p(\Xi,m|\dot{x},\Theta(x))=\frac{p(\Xi,m)p(\dot{x}|\Xi,m,\Theta(x))}{p(\dot{x})}$ (5)
@@
$p(\Xi,m|\dot{x},\mathcal{D})=\frac{p(\Xi,m)p(\mathcal{D}|\Xi,m)}{p(\mathcal{D})}$ (6)
Access to the posterior distribution over both the model terms and its parameterisation confers a number of advantages for the data-driven discovery of nonlinear dynamics. If samples from the posterior are available, a major advantage is that they can be marginalised (in the Monte-Carlo sense) to give direct access to a posterior distribution over models, independently of their parameterisation i.e. $p(m|\mathcal{D})$. This allows the practitioner to evaluate the probability that a term should be included in a model (given the observed data). This permits more robust ways to introduce model parsimony, for example by excluding terms with an inclusion probability below a certain threshold. Alternatively, the full posterior over model terms and parameters can be propogated to further analyses.
@@ 3.2 Reversible-jump MCMC
$\pi(\theta^{\prime})\kappa(\theta|\theta^{\prime})=\pi(\theta)\kappa(\theta^{\prime}|\theta)$ (7)
@@
$k(\theta^{\prime}|\theta)=p(\theta^{\prime}|\theta)\alpha(\theta\rightarrow\theta^{\prime})$ (8)
@@
$\alpha(\theta\rightarrow\theta^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{\pi(\theta)}{\pi(\theta)}}_{\begin{subarray}{c}\text{Target}\\ \text{ratio}\end{subarray}}\underbrace{\frac{p(\theta|\theta^{\prime})}{p(\theta^{\prime}|\theta)}}_{\begin{subarray}{c}\text{Proposal}\\ \text{ratio}\end{subarray}}\Big{\}}$ (9)
@@
In the case that $\theta^{\prime}$ and $\theta$ have different dimensions, detailed balance will not hold for standard choices of the transition density. To overcome this issue, in 1995 Green introduced the reversible-jump MCMC (RJMCMC) algorithm [20] as a method to sample from distributions defined over parameters of different dimension. The RJMCMC method overcomes this difficulty via dimension matching. Let $k$ be the dimension of the state of a Markov chain at $\theta$. In order to move to a new state with dimension $k^{\prime}$ and position $\theta^{\prime}$, an auxiliary variable $u^{\prime}$ with dimension $j^{\prime}$ is sampled. To compensate for the mismatch between dimensions, it is required that $j+k=j^{\prime}+k^{\prime}$, where $j$ is the current dimension of the auxillary variable $u$. Then, a bijective map $g_{k\rightarrow k^{\prime}}:\{u,\theta\}\rightarrow\{u^{\prime},\theta^{\prime}\}$ between each pair of dimensions is defined, such that the dimension of $\{u,\theta\}$ is unchanged by the bijection. If the probability of ‘jumping’ from one dimension to another (via the appropriate corresponding bijection) is given by $J(k^{\prime},\theta^{\prime}|k,\theta)q(u^{\prime}|\theta^{\prime})$ Then the RJMCMC acceptance probability can be reformulated as,
@@
$\alpha(\theta\rightarrow\theta^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{\pi(\theta^{\prime})}{\pi(\theta)}}_{\begin{subarray}{c}\text{Target}\\ \text{ratio}\end{subarray}}\underbrace{\frac{J(k^{\prime},\theta^{\prime}|k,\theta)}{J(k,\theta|k^{\prime},\theta^{\prime})}}_{\text{Jump ratio}}\underbrace{\frac{q(u^{\prime}|\theta^{\prime})}{q(u|\theta)}}_{\begin{subarray}{c}\text{Auxillary}\\ \text{ratio}\end{subarray}}\underbrace{\left|\frac{\partial g_{k\rightarrow k^{\prime}}}{\partial(u,\theta)}\right|}_{\begin{subarray}{c}\text{Jacobian}\\ \text{determinant}\end{subarray}}\Big{\}}$ (10)
@@
In the general setting, working with RJMCMC can be cumbersome for the practitioner. The specification of the bijective maps is non-trivial and has a strong effect on the sampling efficiency of the scheme. Fortunately, a significant simplification is sometimes available and will be employed here. The trick is to sample model parameters independently between models. This is achieved by stetting $u\triangleq\theta^{\prime}$ and $u^{\prime}\triangleq\theta$ in each transition. Thus, each bijection can be defined,
$g_{k\rightarrow k^{\prime}}(u,\theta)=\{u^{\prime},\theta^{\prime}\}=\{\theta,\theta^{\prime}\}$ (11)
@@
Interrogating the above, it is clear that the dimension matching constraint is satisfied and that the Jacobian is a row-wise re-ordering of the identity matrix. Therefore, the Jacobian determinant term in the acceptance ratio must be identically equal to one. The acceptance probability may now be simplified,
@@
$\alpha(\theta\rightarrow\theta^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{\pi(\theta^{\prime})}{\pi(\theta)}}_{\begin{subarray}{c}\text{Target}\\ \text{ratio}\end{subarray}}\underbrace{\frac{J(k^{\prime},\theta^{\prime}|k,\theta)}{J(k,\theta|k^{\prime},\theta^{\prime})}}_{\text{Jump ratio}}\underbrace{\frac{q(\theta|\theta^{\prime})}{q(\theta^{\prime}|\theta)}}_{\begin{subarray}{c}\text{Proposal}\\ \text{ratio}\end{subarray}}\Big{\}}$ (12)
3.3 Bayesian identification of nonlinear dynamics
@@
With the relevant RJMCMC theory established, attention can be returned to the problem of sampling from the posterior density $p(\Xi,m|\mathcal{D})$. In order to develop the posterior over both the model terms and their parameterisation prior and proposal densities must be specified. It is clear that there are many appropriate choices for these objects and in practice it can be expected that prior selection will be guided by the problem at hand. However, it can be shown that certain choices of the parameter proposal can lead to drastic simplification of the inference scheme.
@@
Considering for now the case whereby the noise variance $\sigma^{2}$ in (3) is known and substituting the required posterior distribution into the acceptance probability above, one has,
@@
$\alpha(\Xi\rightarrow\Xi^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{p(\Xi^{\prime},m^{\prime}|\mathcal{D},\sigma^{2})}{p(\Xi,m|\mathcal{D},\sigma^{2})}}_{\begin{subarray}{c}\text{Posterior}\\ \text{ratio}\end{subarray}}\underbrace{\frac{J(m^{\prime}|m)}{J(m|m^{\prime})}}_{\text{Jump ratio}}\underbrace{\frac{q(\Xi|\Xi^{\prime})}{q(\Xi^{\prime}|\Xi)}}_{\begin{subarray}{c}\text{Proposal}\\ \text{ratio}\end{subarray}}\Big{\}}$ (13)
@@
The proposal density in the above is freely chosen by the user. A natural choice, (after Troughton and Godsill [22]) is to employ the full conditional density,
$q(\Xi^{\prime}|\Xi)=p(\Xi^{\prime}|m^{\prime},\mathcal{D},\sigma^{2})$ (14)
substituting into the above, one finds that since,
$p(m|\mathcal{D},\sigma^{2})=\frac{p(\Xi,m|\mathcal{D},\sigma^{2})}{p(\Xi|m,\mathcal{D},\sigma^{2})}$ (15)
@@
$\alpha(\Xi\rightarrow\Xi^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{p(m^{\prime}|\mathcal{D},\sigma^{2})}{p(m|\mathcal{D},\sigma^{2})}}_{\begin{subarray}{c}\text{Ratio of model}\\ \text{posteriors}\end{subarray}}\underbrace{\frac{J(m^{\prime}|m)}{J(m|m^{\prime})}}_{\text{Jump ratio}}\Big{\}}$ (16)
This simplification is only available for particular choices of the prior $p(\Xi,m)$ that are conjugate, e.g. the Gaussian likelihood defined above (and defined independently to the prior over models). However, it gives direct access to the marginal posteriors $p(m|\mathcal{D})$ (see [23, 24] for example). Considering here a conjugate Gaussian prior over the parameter values (independently of their inclusion in the model), one has,
$p(\Xi|m)\triangleq\mathcal{N}(\mu_{m}^{(0)},\Sigma_{m}^{(0)})$ (17)
$\begin{split}p(m|\mathcal{D},\sigma^{2})&\propto p(m)p(\mathcal{D},\sigma^{2}|m,\Xi)\\ &\propto p(m)\sigma^{-k}|\Sigma_{m}|^{\frac{1}{2}}\exp\left(\frac{1}{2}\mu_{m}^{\mkern-1.5mu\mathsf{T}}\Sigma^{-1}_{m}\mu_{m}\right)\end{split}$ (18)
@@
$\Sigma_{m}=\sigma^{2}\left[\Theta_{m}^{\mkern-1.5mu\mathsf{T}}\Theta_{m}+\Sigma_{m}^{(0)-1}\right]^{-1}$ (19)
$\mu_{m}=\Sigma_{m}\Sigma_{m}^{(0)-1}\mu_{m}^{(0)}+\frac{1}{\sigma^{2}}\Sigma_{m}^{\mkern-1.5mu\mathsf{T}}\Theta^{\mkern-1.5mu\mathsf{T}}\dot{x}$ (20)
are the posterior mean and covariance of the parameters for a given model $m$ (see [23, 24] for example). Thus, the overall acceptance probability of a move from $\Xi$ to $\Xi^{\prime}$ can be written,
$\alpha(\Xi\rightarrow\Xi^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{p(m^{\prime})}{p(m)}}_{\text{Model prior}}\underbrace{\frac{J(m^{\prime}|m)}{J(m|m^{\prime})}}_{\text{Jump ratio}}\underbrace{\frac{\sigma^{-k^{\prime}}|\Sigma_{m^{\prime}}|^{\frac{1}{2}}\exp\left(\frac{1}{2}\mu_{m^{\prime}}^{\mkern-1.5mu\mathsf{T}}\Sigma^{-1}_{m^{\prime}}\mu_{m^{\prime}}\right)}{\sigma^{-k}|\Sigma_{m}|^{\frac{1}{2}}\exp\left(\frac{1}{2}\mu_{m}^{\mkern-1.5mu\mathsf{T}}\Sigma^{-1}_{m}\mu_{m}\right)}}_{\begin{subarray}{c}\text{Model posterior ratio}\end{subarray}}\Big{\}}$ (21)
The above can now be readily used to sample from the required posterior. In practice, the noise variance in (3) will be unknown. This problem can be simply addressed by including a Gibbs move for the noise variance given the model and the parameters, which are available following a reversible-jump move. In this work, an inverse-gamma prior is used. The Gibbs step is therefore,
$p(\sigma^{2})\triangleq\mathcal{IG}(a^{(0)},b^{(0)})$ (22)
@@
$p(\sigma^{2}|\Xi,m,\mathcal{D})=\mathcal{IG}(a,b)$ (23)

### v2

@@ 1 Introduction
Bayesian uncertainty quantification is a natural framework for conducting UQ in data-driven modelling. Indeed, much has been written on the topic, e.g. [2, 3]. Physical knowledge and inductive biases (such as parsimony) can be incorporated as prior knowledge, making such modelling assumptions explicit. Indeed, many literature contributions propose Bayesian methods for UQ in both model parameterisation [4, 5, 6] and model structure [7, 8, 9, 10]. Further discussion of specific related works is presented in Section 3.2.
@@ 1.1 Contribution
The remainder of this paper is structured as follows: The following section introduces the necessary background, summarising the SINDy methodology, holding some discussion on parsimony in data-driven modelling motivating the use of reversible-jump Monte-Carlo. A third section introduces the proposed approach and its position in the literature, providing a framework for Bayesian identification of nonlinear dynamics. A fourth section compares the proposed approach to the ensemble SINDy method in three benchmark case studies. A final section presents some discussion and directions for future work.
2 Background
2.1 SINDy methods
The SINDy method proposed, in 2016 by Brunton et al. [1], has received considerable attention in the scientific literature as a computationally-efficient way to learn differential equations directly from data. Since it was first proposed, the method has been proven to be effective in correctly identifying governing equations from both simulated and real-world data.
At their core, SINDy methods make the key assumption that the observed dynamics admit a first-order formulation that is linear in the parameters of a library of user-specified basis functions $\Theta$. Let $f$ denote the first-order ordinary differential equation (ODE) that describes the evolution of some state-space dynamics $x(t)$, [^1] The method can be extended to systems of ODEs, PDEs etc. but for notational simplicity, the exposition in this paper will consider scalar valued ODEs. The authors note that the extension to systems of ODEs (as seen in the case-studies of this manuscript) can be handled by treating each of the regression targets $\dot{x}_{i}(t)$ independently, from a shared state space comprised of all $x_{i}$,
$\dot{x}(t)=f(x(t)){\color[rgb]{0,0,0},}$ (1)
@@ 2.1 SINDy methods
$\dot{x}\approx\Theta(x)\Xi{\color[rgb]{0,0,0},}$ (2)
where $x$ is the measured state, $\Theta(x)\in\mathbb{R}^{N\times D}$ is a library of $D$ user-selected basis functions computed from $x$. In the case that $x$ is a vector, the library of terms may also include functions of all elements in $x$ and cross terms. Thus, it is expected that $D$ will grow much faster than the dimension of the ODE under investigation. Critically, the coefficient vector $\Xi$ is assumed to be sparse, whereby many elements of $\Xi$ are equal to 0. In practice, it is not often the case that measurements of both $x$ and $\dot{x}$ are available. It is usual in the SINDy literature to assume that only the states, $x$, are measured, and their derivatives $\dot{x}$ must be computed numerically (with sufficient accuracy such that (2) holds) [1]. The SINDy framework (for identification of ODEs) can thus be summarised by three steps:
@@
Of the three components in SINDy methods described here, this paper is concerned only with the third—determination of appropriate model terms and their parameters. At this stage, it is useful to present some discussion as to the role of parsimony in data-driven modelling.
@@ 2.2 Parsimony in data-driven modelling
A particular issue with parameter size as a proxy for term importance is the selection of a threshold value, below which model terms should be excluded. In practice, this level can be difficult to select in a principled manner, leading to a spectrum of models at different levels of sparsity. In [20], an approach based on the Akaike information criterion is used to establish a Pareto front of models trading off error against the number of identified terms. This approach is successful in that the user is offered an interpretable choice from a small subset of possible models. However, the theoretical interpretation of this Pareto front is limited and distributional estimates are not explicitly available. The method also relies on parameter size as a proxy for term importance which can lead to the issues described above.
@@
${\color[rgb]{0,0,0}\dot{x}=\Theta(x)\Xi+\epsilon,\quad\epsilon\sim\mathcal{N}(0,\sigma^{2}),}$ (3)
where $\sigma^{2}$ is the noise variance in the observation model. The model identification task at hand is to identify a subset of the columns of $\Theta(x)$. Intuitively, this model set includes all subsets of library functions (the powerset) and thus all possible ways that the parameter vector may be sparsified. Let
=$\mathcal{M}=\{m_{i}\}_{i=1}^{2^{n}}{\color[rgb]{0,0,0},}$ (4)
@@
$p(\Xi,m|\dot{x},\Theta(x))=\frac{p(\Xi,m)p(\dot{x}|\Xi,m,\Theta(x))}{p(\dot{x})}{\color[rgb]{0,0,0}.}$ (5)
@@
$p(\Xi,m|\dot{x},\mathcal{D})=\frac{p(\Xi,m)p(\mathcal{D}|\Xi,m)}{p(\mathcal{D})}{\color[rgb]{0,0,0}.}$ (6)
Access to the posterior distribution over both the model terms and its parameterisation confers a number of advantages for the data-driven discovery of nonlinear dynamics. If samples from the posterior are available, a major advantage is that they can be marginalised (in the Monte-Carlo sense) to give direct access to a posterior distribution over models, independently of their parameterisation i.e. $p(m|\mathcal{D})$. This allows the practitioner to evaluate the probability that a term should be included in a model (given the observed data). This permits more robust ways to introduce model parsimony, for example by excluding terms with an inclusion probability below a certain threshold. Alternatively, the full posterior over model terms and parameters can be propagated to further analyses.
@@ 2.3 Reversible-jump MCMC
$\pi(\theta^{\prime})\kappa(\theta|\theta^{\prime})=\pi(\theta)\kappa(\theta^{\prime}|\theta){\color[rgb]{0,0,0}.}$ (7)
@@
$k(\theta^{\prime}|\theta)=p(\theta^{\prime}|\theta)\alpha(\theta\rightarrow\theta^{\prime}){\color[rgb]{0,0,0},}$ (8)
@@
$\alpha(\theta\rightarrow\theta^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{\pi(\theta)}{\pi(\theta)}}_{\begin{subarray}{c}\text{Target}\\ \text{ratio}\end{subarray}}\underbrace{\frac{p(\theta|\theta^{\prime})}{p(\theta^{\prime}|\theta)}}_{\begin{subarray}{c}\text{Proposal}\\ \text{ratio}\end{subarray}}\Big{\}}{\color[rgb]{0,0,0},}$ (9)
@@
In the case that $\theta^{\prime}$ and $\theta$ have different dimensions, detailed balance will not hold for standard choices of the transition density. To overcome this issue, in 1995 Green introduced the reversible-jump MCMC (RJMCMC) algorithm [21] as a method to sample from distributions defined over parameters of different dimension. The RJMCMC method overcomes this difficulty via dimension matching. Let $k$ be the dimension of the state of a Markov chain at $\theta$. In order to move to a new state with dimension $k^{\prime}$ and position $\theta^{\prime}$, an auxiliary variable $u^{\prime}$ with dimension $j^{\prime}$ is sampled. To compensate for the mismatch between dimensions, it is required that $j+k=j^{\prime}+k^{\prime}$, where $j$ is the current dimension of the auxiliary variable $u$. Then, a bijective map $g_{k\rightarrow k^{\prime}}:\{u,\theta\}\rightarrow\{u^{\prime},\theta^{\prime}\}$ between each pair of dimensions is defined, such that the dimension of $\{u,\theta\}$ is unchanged by the bijection. If the probability of ‘jumping’ from one dimension to another (via the appropriate corresponding bijection) is given by $J(k^{\prime},\theta^{\prime}|k,\theta)q(u^{\prime}|\theta^{\prime})$ Then the RJMCMC acceptance probability can be reformulated as,
@@
$\alpha(\theta\rightarrow\theta^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{\pi(\theta^{\prime})}{\pi(\theta)}}_{\begin{subarray}{c}\text{Target}\\ \text{ratio}\end{subarray}}\underbrace{\frac{J(k^{\prime},\theta^{\prime}|k,\theta)}{J(k,\theta|k^{\prime},\theta^{\prime})}}_{\text{Jump ratio}}\underbrace{\frac{q(u^{\prime}|\theta^{\prime})}{q(u|\theta)}}_{\begin{subarray}{c}\text{Auxillary}\\ \text{ratio}\end{subarray}}\underbrace{\left|\frac{\partial g_{k\rightarrow k^{\prime}}}{\partial(u,\theta)}\right|}_{\begin{subarray}{c}\text{Jacobian}\\ \text{determinant}\end{subarray}}\Big{\}}{\color[rgb]{0,0,0}.}$ (10)
@@
In the general setting, working with RJMCMC can be cumbersome for the practitioner. The specification of the bijective maps is non-trivial and has a strong effect on the sampling efficiency of the scheme. Fortunately, a significant simplification is sometimes available and will be employed here. The trick is to sample model parameters independently between models. This is achieved by setting $u\triangleq\theta^{\prime}$ and $u^{\prime}\triangleq\theta$ in each transition. Thus, each bijection can be defined, as ‘exchanging’ the auxiliary parameters with new ones in the model move,
$g_{k\rightarrow k^{\prime}}(u,\theta)=\{u^{\prime},\theta^{\prime}\}=\{\theta,\theta^{\prime}\}{\color[rgb]{0,0,0}.}$ (11)
@@
Considering the above, it is clear that the dimension matching constraint is satisfied and that the Jacobian is a row-wise re-ordering of the identity matrix. Therefore, the Jacobian determinant term in the acceptance ratio must be identically equal to one. The acceptance probability may now be simplified to
@@
$\alpha(\theta\rightarrow\theta^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{\pi(\theta^{\prime})}{\pi(\theta)}}_{\begin{subarray}{c}\text{Target}\\ \text{ratio}\end{subarray}}\underbrace{\frac{J(k^{\prime},\theta^{\prime}|k,\theta)}{J(k,\theta|k^{\prime},\theta^{\prime})}}_{\text{Jump ratio}}\underbrace{\frac{q(\theta|\theta^{\prime})}{q(\theta^{\prime}|\theta)}}_{\begin{subarray}{c}\text{Proposal}\\ \text{ratio}\end{subarray}}\Big{\}}{\color[rgb]{0,0,0}.}$ (12)
3 Bayesian identification of nonlinear dynamics
@@
With the relevant RJMCMC theory established, and the inference problem formalised, attention can be returned to the problem of sampling from the posterior density $p(\Xi,m|\mathcal{D})$. Development of this posterior over both the model terms, and their parameterisation requires the specification of both prior and proposal densities. It is clear that there are many appropriate choices for these objects and, in practice, it can be expected that prior selection will be guided by the problem at hand. However, it can be shown that certain choices of the parameter proposal can lead to drastic simplification of the inference scheme.
@@
Considering for now the case whereby the noise variance $\sigma^{2}$ in (3) is known and by substituting the required posterior distribution into the acceptance probability (12) above, one has,
@@
$\alpha(\Xi\rightarrow\Xi^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{p(\Xi^{\prime},m^{\prime}|\mathcal{D},\sigma^{2})}{p(\Xi,m|\mathcal{D},\sigma^{2})}}_{\begin{subarray}{c}\text{Posterior}\\ \text{ratio}\end{subarray}}\underbrace{\frac{J(m^{\prime}|m)}{J(m|m^{\prime})}}_{\text{Jump ratio}}\underbrace{\frac{q(\Xi|\Xi^{\prime})}{q(\Xi^{\prime}|\Xi)}}_{\begin{subarray}{c}\text{Proposal}\\ \text{ratio}\end{subarray}}\Big{\}}{\color[rgb]{0,0,0}.}$ (13)
@@
The proposal density $q$ in (13) is freely chosen by the user. A natural choice, (after Troughton and Godsill [23] and similar to the approach of [24]) is to employ the full conditional density,
$q(\Xi^{\prime}|\Xi)=p(\Xi^{\prime}|m^{\prime},\mathcal{D},\sigma^{2}){\color[rgb]{0,0,0}.}$ (14)
Substituting proposal into (13) above, one finds that since,
$p(m|\mathcal{D},\sigma^{2})=\frac{p(\Xi,m|\mathcal{D},\sigma^{2})}{p(\Xi|m,\mathcal{D},\sigma^{2})}{\color[rgb]{0,0,0},}$ (15)
@@
$\alpha(\Xi\rightarrow\Xi^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{p(m^{\prime}|\mathcal{D},\sigma^{2})}{p(m|\mathcal{D},\sigma^{2})}}_{\begin{subarray}{c}\text{Ratio of model}\\ \text{posteriors}\end{subarray}}\underbrace{\frac{J(m^{\prime}|m)}{J(m|m^{\prime})}}_{\text{Jump ratio}}\Big{\}}{\color[rgb]{0,0,0}.}$ (16)
This simplification is only available for particular choices of the prior $p(\Xi,m)$ that are conjugate, e.g. the Gaussian likelihood defined above (and defined independently to the prior over models). Moreover, it gives direct access to the marginal posteriors $p(m|\mathcal{D})$ (see [25] (Chapter 6) or [26] (Chapter 3) for example). Considering here, a conjugate Gaussian prior at each model order (conditioned on the model choice), one has,
$p(\Xi|m)\triangleq\mathcal{N}(\mu_{m}^{(0)},\Sigma_{m}^{(0)}){\color[rgb]{0,0,0}.}$ (17)
Because the prior is conjugate, for a given model and noise variance, one may write,
${\color[rgb]{0,0,0}p(\Xi|m,\mathcal{D},\sigma^{2})=\mathcal{N}(\mu_{m},\Sigma_{m})}{\color[rgb]{0,0,0},}$ (18)
@@
$\Sigma_{m}=\sigma^{2}\left[\Theta_{m}^{\mkern-1.5mu\mathsf{T}}\Theta_{m}+\Sigma_{m}^{(0)-1}\right]^{-1}{\color[rgb]{0,0,0},}$ (19)
$\mu_{m}=\Sigma_{m}\Sigma_{m}^{(0)-1}\mu_{m}^{(0)}+\frac{1}{\sigma^{2}}\Sigma_{m}^{\mkern-1.5mu\mathsf{T}}\Theta^{\mkern-1.5mu\mathsf{T}}\dot{x}{\color[rgb]{0,0,0},}$ (20)
are the posterior mean and covariance of the parameters for a given model $m$ (see [25, 26] for example). The marginal posteriors are thus available as,
$\begin{split}p(m|\mathcal{D},\sigma^{2})&\propto p(m)p(\mathcal{D},\sigma^{2}|m,\Xi){\color[rgb]{0,0,0},}\\ &\propto p(m)\sigma^{-k}|\Sigma_{m}|^{\frac{1}{2}}\exp\left(\frac{1}{2}\mu_{m}^{\mkern-1.5mu\mathsf{T}}\Sigma^{-1}_{m}\mu_{m}\right){\color[rgb]{0,0,0}.}\end{split}$ (21)
Thus, the overall acceptance probability of a move from $\Xi$ to $\Xi^{\prime}$ can be written,
$\alpha(\Xi\rightarrow\Xi^{\prime})=\min\Big{\{}1,\\ \underbrace{\frac{p(m^{\prime})}{p(m)}}_{\text{Model prior}}\underbrace{\frac{J(m^{\prime}|m)}{J(m|m^{\prime})}}_{\text{Jump ratio}}\underbrace{\frac{\sigma^{-k^{\prime}}|\Sigma_{m^{\prime}}|^{\frac{1}{2}}\exp\left(\frac{1}{2}\mu_{m^{\prime}}^{\mkern-1.5mu\mathsf{T}}\Sigma^{-1}_{m^{\prime}}\mu_{m^{\prime}}\right)}{\sigma^{-k}|\Sigma_{m}|^{\frac{1}{2}}\exp\left(\frac{1}{2}\mu_{m}^{\mkern-1.5mu\mathsf{T}}\Sigma^{-1}_{m}\mu_{m}\right)}}_{\begin{subarray}{c}\text{Model posterior ratio}\end{subarray}}\Big{\}}{\color[rgb]{0,0,0}.}$ (22)
The acceptance probability in (22) can now be readily used to sample from the required posterior. In practice, the noise variance in (3) will be unknown. This problem can be simply addressed by including a Gibbs move for the noise variance given the model and the parameters, which are available following a reversible-jump move. For example an inverse-gamma prior, conjugate to the Gaussian likelihood can be employed. The Gibbs step in this case is,
$p(\sigma^{2})\triangleq\mathcal{IG}(a^{(0)},b^{(0)}){\color[rgb]{0,0,0}.}$ (23)
@@
$p(\sigma^{2}|\Xi,m,\mathcal{D})=\mathcal{IG}(a,b){\color[rgb]{0,0,0},}$ (24)